David Fifield’s Eagle Project

In order for a Boy Scout to achieve the rank of Eagle, he must demonstrate his leadership ability in a project that benefits the community. For my project, I have chosen to collect books for the Aurora Public Library. I am currently accepting donations of used books or money to purchase new books. Used books will either be sold through the library’s used book store to generate funds, or placed in the library collection, depending on the library’s needs. Books of all kinds are welcome, but I have chosen to concentrate on collecting books dealing with computers and technology, as these are personal interests of mine. My goal is to make the library’s technology section both current and complete.

These are several books which I feel are particularly valuable:

• Any of Donald E. Knuth’s excellent series of books The Art of Computer Programming, published by Addison Wesley (http://www.aw.com/).

• The C Programming Language, Second Edition by Brian W. Kernighan and Dennis M. Ritchie, published by Prentice-Hall (http://www.phptr.com/).

• O’Reilly and Associates (http://www.ora.com/) publishes many good books on varied computer-related topics, each identifiable by an animal on the cover.

Some fields change so fast that the books that discuss them are updated frequently. For this reason, I am also accepting donations of money to purchase the latest editions of these books.

I have attempted to make contacting me regarding this project as easy as possible. This

project has a web site at http://171recon.freeservers.com/eagle/. This web site is

the source of the most up-to-date information regarding this project, including a hypertext version of this flyer (perfect for emailing to family and friends). Additionally, I may be contacted by email at david@171recon.freeservers.com with questions or to arrange a donation. If you are unable to reach me by either of these two preferred methods, you may call me at (303) 766–3525.

I’d like to thank you in advance for your support in this endeavor. I’m certain you will be able to appreciate the many benefits that more books available to local youth will bring.

Yours in Scouting,

David Fifield

Boy Scout Troop 171